OpenStudy (anonymous):

If BC= 4 and CE=15 write 2 lengths for segments AC and CD are: HELPPPPPPPPPP

OpenStudy (anonymous):



OpenStudy (anonymous):

are the lines intersecting the center? do they touch the center?

OpenStudy (anonymous):

No none touch the center

OpenStudy (anonymous):

how is bc=4?

OpenStudy (anonymous):

IDK its hard

OpenStudy (mathstudent55):

BC*CE = AC*CD

OpenStudy (anonymous):

Huuhhh...

OpenStudy (anonymous):

Do both lines go through the center?

OpenStudy (anonymous):

@Rancd

OpenStudy (mathstudent55):

The product of the lengths of the segments of one chord equals the product of the lengths of the segments of the other chord.

jimthompson5910 (jim_thompson5910):

BC*CE = AC*CD 4*15 = AC*CD 60 = AC*CD AC*CD = 60 So if AC and CD are whole numbers, then just list out the different ways to factor 60

OpenStudy (anonymous):

Factors of 60:?

OpenStudy (anonymous):

5, 10

OpenStudy (mathstudent55):

Since you know that BC = 4 and CE = 15, their product is 4*15 = 60 That means the product of AC and CD must also equal 60. Pick any two numbers for AC and CD whose product is 60.

OpenStudy (anonymous):

Rancd, do both lines go through the center????????????????????

OpenStudy (mathstudent55):

You mean 6, 10

OpenStudy (anonymous):

Yeah and no thy do not intersect at the center

OpenStudy (anonymous):

Hey I'm official confused

jimthompson5910 (jim_thompson5910):

6*10 = 60, so AC could be 6 and CD could be 10 and this would work out so AC = 6 CD = 10 or you can flip them around to get the same thing

OpenStudy (anonymous):

I get it noww

jimthompson5910 (jim_thompson5910):

you can use 1*60, 2*30, 3*20, 5*12 or flip either of them around

OpenStudy (anonymous):

Thanks guys you help me with my test review now I'll definitly ace it